OpenStudy (gabylovesyou):

Algebra 1 help square root.. :p

OpenStudy (anonymous):

yah sure tll

OpenStudy (anonymous):

What square roots do you need help with?? Jw.

OpenStudy (gabylovesyou):

6sqrt(v) - sqrt(4v^2) - sqrt(36v) + sqrt(v^2)

OpenStudy (anonymous):

What grade are you in?

OpenStudy (gabylovesyou):

8th ... but im advanced so 9th grade math

OpenStudy (anonymous):

- sqrt(4v^2) = - sqrt[ (2v)^2 ] = - |2v| do you need help on the others too?

OpenStudy (gabylovesyou):

yes its an equation... adding and subtracting them..... how do i do that ^

OpenStudy (anonymous):

let me try to show you the steps. Anything you dont understand just ask

OpenStudy (gabylovesyou):

ok thanks :)

OpenStudy (anonymous):

6sqrt(v) - sqrt(4v^2) - sqrt(36v) + sqrt(v^2) = 6sqrt(v) - sqrt[ (2v)^2 ] - sqrt[ v*6^2] + |v| = 6*v^(1/2) - |2v| - 6*v^(1/2) + |v| is this better for you?

OpenStudy (anonymous):

well that's your teacher resolution?

OpenStudy (anonymous):

I mean did you learn this from your teacher? sqrt(v^2)=v; sqrt(4v^2)=sqrt(4)*sqrt(v^2)=2v; This 2 particular cases?

OpenStudy (anonymous):

Yes.

OpenStudy (gabylovesyou):

i get -v :)

OpenStudy (anonymous):

well i don't want to tell you that she is a bad teacher or something because i don't know her but when you have a square variable inside a square root you cant simply cut the 2 exponents.... You have to use the modulus sign as the variable could be negative aswell

OpenStudy (gabylovesyou):

is -v correct ?

OpenStudy (anonymous):

Okay then. You answer the question and i'll delete my answer.

OpenStudy (anonymous):

example: sqrt(v^2) if v = -2 then: sqrt( (-2)^2) = sqrt(4) = 2 if v = 2 then: sqrt( (2)^2) = sqrt(4) = 2 as you can see both -2 and 2 are a solution. Thats why when you cut the exponents you should use modulus signs The correct answer would be sqrt(v^2) = |v|

OpenStudy (anonymous):

@Gabylovesyou you now should calculate that equations for the cases that |v| = -v and for |v| = v

OpenStudy (gabylovesyou):

v is not in my choices.....

OpenStudy (gabylovesyou):

3v v sqrt(v) -v 6v^2

OpenStudy (anonymous):

-v is the correct answer. You are correct :)

OpenStudy (gabylovesyou):

ok thanks :)

OpenStudy (anonymous):

yaaay. Gabylovesme

OpenStudy (gabylovesyou):

lol :p